OpenStudy (anonymous):

Find the average rate of change of the function f over the given interval. f(x)=x^2-6/4x-8 from x=6 to x=8

OpenStudy (anonymous):

\[\frac{f(8)-f(6)}{8-6}\] \[\frac{f(8)-f(6)}{2}\] they should be easy enough to find i hope

OpenStudy (anonymous):

Okay, so would you plug those numbers in and find the answer that way? I am so lost and confused. Fail.

OpenStudy (anonymous):

you need two number now, \[f(8)\] and \[f(6)\] is it clear how to find them?

OpenStudy (anonymous):

Not really, can you explain step by step?

OpenStudy (anonymous):

\[f(x)=\frac{x^2-6}{4x-8}\] \[f(8)=\frac{8^2-6}{4\times 8-8}\]

OpenStudy (anonymous):

so that would be 29/12 right?